OpenStudy (jennychan12):

find lim x-->0 (sin(3x))/x

OpenStudy (jennychan12):

\[\lim_{x \rightarrow 0} \frac{ \sin(3x) }{ x }\]

OpenStudy (campbell_st):

this need to be rewritten and use the following limit fact \[\lim_{\theta \rightarrow 0} \frac{\sin(\theta)}{\theta} = 1\] so rewriting the fraction so the denominator is the same as the angle it can be written as \[\lim_{x \rightarrow 0}\frac{3\sin(3x)}{3x}\] I let h = 3x in the limit below... \[\ = 3 \times \lim_{h \rightarrow 0}\frac{\sin(h)}{h} = 3 \times 1\]

OpenStudy (jennychan12):

ok thanks.